Systems Staff Meeting Notes 6/25/01

Systems Staff Meeting Notes
June 25, 2001

Present: Cynthia Bowker, Will Chaney, Bill Covey, Debra Harris, Winston Harris, Suzy Shaw, Hilton Soto, Geoffrey Stewart, Ying Tang
Announcements
  1. Bill and Suzy were married in Wisconsin last week! A website with pictures of the event will be forthcoming.
  2. Ying is the Florida Women’s Single Champion of the Sunshine State Summer Games in badminton!
  3. Cox Cable will have high-speed cable access available by the end of this month. It is the most cost-effective for fast access – but it does operate on a party line. This could leave you vulnerable to attack, might want to invest in a baby firewall.
  4. The Tape Vault (new server room) door has recently been fobbed. No access possible without a properly programmed key fob or card. All of Systems should have access.
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. Compaq Web Server – problems with the company regarding this order – they say that they are unable to find the RAID controller. Bill talked to the sales rep, she said that they have shipped pieces of it. Bill told her to make sure to have a technician available to assemble it & make sure it works.
  2. Bill notes rumors that the governor has decreed that there will be no carry-forward funds. If this is accurate, all orders will need to be received and paid for by June 30.
  3. Color LaserJet printer for Judaica – pending since April. Debra talked with Laser Action Plus – faxed them the PO – it should be arriving this week.
  4. We will continue to process orders as we receive them. If they have to hold them in Accounting, that’s okay.
  5. Jim Arnette left a voicemail message with Bill regarding the pending money that is scheduled to expire on our NERDC mainframe account. Bill wants everyone to know that all the money is in the first ordinal of account #001, the funds are for the library account 10095001 & the ID is SYSMAST. Funding should all be distributed to the central account. We distribute it out when we run our funding program. Library Accounting sends an encumbrance to Central Finance & Accounting, but we usually end up having to send a copy directly to NERDC to receive the funds. Often, come July 1st, the funds may not be in place yet and we will get calls regarding Dean’s Menu problems.
  6. Problems with NOTIS dropping out this morning. Could be network problems in West - 3270 sessions are being affected. Problems are ghostly in nature – not repeatable. Could be due to changes made by NERDC over the weekend.
Networking
  1. Big monster server, SmathersNT2R, is finally recognizing the backup exec tapes. Will hopes to get departmental folders set up for storage purposes. The mail server will eventually be on it as well. Org chart will be necessary – will probably be a good idea to upgrade to W2K before the Exchange server move as well.
  2. Remote installation of W2K has been hit and miss. Geoffrey has been working on burning images on CD for remote installation. Some remote installs have been a fiasco due to unique software or partition problems. If we do fresh installs instead of upgrades, it should go more successfully. We can do the installation from the server, but problems do exist. We could try to do the upgrades first then a fresh install if we run into problems. The CD images can be accessed across the network if the file is on the local subnet server.
  3. Squid – trouble last week – tried to reboot on Thursday, had to remove the cache. Restarted after that and it seems to be fixed. We seem to be overrunning the cache inside – will try to tune the configuration file.
  4. MSL – reference desk is to be removed today – they need wires cut in order to do so.
Software Projects Status
  1. Trouble Tickets are now multi-user.
  2. Jack Waters requested some Library Statistics mods.
  3. DLC – moved folders from Disk 1 to Disk 2 for Erich – freed up some much-needed disk space.
  4. Hilton – had an efficiency request from DLC to move data into a database. It is the Perl program written by Maureen – he’s transferring it to Delphi.
  5. IRDB changes made.
  6. Diana Hagan – Hilton worked on a Java Script project for her – geographic coordinates for MARC records.
  7. Ying finished the Sniff interface for Richard Phillips. She wrote him a custom view to make it more accurate.
  8. Public workstation authentication – moving slowly, Winston has been communicating with Tigert. He had to create backup groups with production and test regions in CICS.
DP Coordinator
  1. Students are out this week – intercession break.
  2. Student schedules for Summer B are still pending, may still be figuring it out next week.


Back to the Systems Homepage

Last updated July 2, 2001
by Debra Harris
debharr@mail.uflib.ufl.edu